\begin{thebibliography}{99}
 		
 	\bibitem{bass} Bass, H., Connell, E. H., and Wright, D. The Jacobian conjecture: reduction of degree and formal expansion of the inverse. {\em Bulletin of the American Mathematical Society}, 7(2), 287--330 (1982). 
 	
 		
 	\bibitem{Belov_etc} Belov, A., Bokut, L., Rowen, L., Yu, J.-T. 
 	The Jacobian conjecture, together with Specht and Burnside-type problems. In: 
 	James McKernan, Ciro Ciliberto, Mikhail Zaidenberg, Ivan Cheltsov ed., 
 	 {\em Automorphisms in Birational and Affine Geometry.
 	Springer Proceedings in
 	Mathematics and Statistics}, Vol. 79. Cham: Springer, pp. 249--285.  (2014). ArXiv 
 	preprint arXiv:1308.0674.
 	
\bibitem{dots} Dotsenko, V. Nilpotence, weak nilpotence, and the nil property in the nonassociative world: computations and conjectures.  {\em Experimental Mathematics}, (2024). 1-9.  ArXiv 
 	preprint arXiv:2306.11362.
 		
\bibitem{dr} L.M. Dru\. zkowski, K. Rusek, The formal inverse and the Jacobian conjecture.  {\em Ann. Polon. Math.}, 46, 85--90 (1985).
 
 	\bibitem{Gerst1960} Gerstenhaber, M.  On nilalgebras and linear varieties of nilpotent matrices. II.  {\em Duke Mathematical Journal}, 27 (1), 21--31 (1960).
 	

\bibitem{gz} Gorni, G., Zampieri, G. Yagzhev polynomial mappings: on the structure of the Taylor expansion of their local inverse.  {\em Annales Polonici Mathematici}, 64, 3, 285--290, (1996). 

\bibitem{pz} Piontkovski, D., Zitan, F. Bernstein algebras that are algebraic and the Kurosh problem.  {\em Journal of Algebra}, 617, 275--316 (2023). 

\bibitem{singer} Singer, D. On Catalan trees and the Jacobian conjecture.  {\em The electronic journal of combinatorics}, vol. 8, 1, R2, 1--35  (2001).

 		\bibitem{Umirbaev} Umirbaev, U.  Polarization algebras and their relations.  {\em Journal of Commutative Algebra}, 11 (3), 433--451 (2019).
 	
 		\bibitem{Yagzhev} A.V. Yagzhev, Keller’s problem,  {\em Sibirsk. Mat. Zh.}, 21:5 (1980), 141--150; 
 {\em Siberian Math. J.}, 21:5, 747--754 (1980).

\bibitem{Yagzhev2000} A.V. Yagzhev, Endomorphisms of polynomial rings and free algebras of different varieties. (Russian)
 {\em Proceedings of Shafarevich Seminar}, Moscow, 1999--2000, issue 2, 15--47.
 		
 			\bibitem{Wang} Wang, S. S. S.  A Jacobian criterion for separability.  {\em Journal of Algebra}, 65 (2), 453--494 (1980).
 		
 	
 	
\end{thebibliography}

\documentclass{amsart}
\usepackage{latexsym,amsxtra,amscd,ifthen}
\usepackage{amsfonts}
\usepackage{verbatim}
\usepackage{amsmath}
\usepackage{amsthm}
\usepackage{amssymb}

\numberwithin{equation}{section}

\usepackage{amssymb,amsmath}
\usepackage{xcolor}

\newtheorem{Theorem}{Theorem}

\newtheorem{Proposition}[Theorem]{Proposition}

\newtheorem{Conjecture}[Theorem]{Conjecture}

\newcommand{\kk}{{\Bbbk}} 

\newcommand{\Z}{{\mathbb Z}}

\newcommand{\Complex}{{\mathbb C}}

\def\ep{\hfill{\vbox to 7pt{\hbox to 7pt{\vrule height 7pt width 7pt}}}}

\DeclareMathOperator{\AdOper}{Ad}

\newcommand{\Identity}{Id}

\begin{document}
	
	\title{Multilinear nilalgebras and the Jacobian theorem}

	\author{Dmitri Piontkovski}
	
	\address{20 Myasnitskaya ul., National Research University Higher School of Economics, Moscow 101000, Russia
	}
	
	\email{dpiontkovski@hse.ru}
	
	\subjclass[2020]{17A42; 15A39; 14R15}
	

	\date{\today}
	
  \maketitle   	
 
\begin{abstract}
If a symmetric multilinear algebra is weakly nil, then it is Engel.  This result may be regarded as an infinite-dimensional analogue of the well-known Jacobian theorem, which states that if a polynomial mapping has a polynomial inverse, then its Jacobian matrix is invertible. This refines a theorem of Gerstenhaber and partially answers a question posed by Dotsenko. 
\end{abstract}

 \label{sec:Jacobian}

 There are various versions of nil algebras in non-associative and multilinear settings. These notions are closely connected with problems concerning polynomial automorphisms.  

 Let $A$ be a vector space and $\mu: A^d \to A$ be a $d$-linear map, so that the pair $(A,\mu)$ forms a (multilinear) algebra (in the sense of general algebra).
The algebra $A$ is called {\em $n$-Engel} (or simply Engel) if for each element $x\in A$ the linear operator of multiplication by $x$ is $n$-nilpotent, that is, the $n$-th Engel identity 
$$
   \AdOper_x^n(y) \equiv 0 
$$ 
holds in $A$, where $\AdOper^1_x (y)  = \AdOper_x (y) = \mu (x, \dots, x, y), 
 \AdOper^{k+1}_x (y) = \mu (x, \dots, x, \AdOper^{k}_x (y))$. 

Gerstenhaber  connected the Engel property with another kind of nilpotence. We call $A$ {\em Gerstenhaber nil} (of nilindex $n$) if for each element $x\in A$, all multilinear multiple compositions of at least $n$ copies of $\mu$ applied to $x$ are zero.  In other words, $A$ is Gerstenhaber nil if each subalgebra generated by a single element is nilpotent of bounded degree. 

From now on, assume that   the  ground field $\kk$ has zero characteristic and the algebra $A$ is commutative in the sense that the operation $\mu$ is symmetric, i.~e., for every $\sigma \in S_d$ we have
 	$\mu(x_1, \dots , x_d ) = \mu(x_{\sigma 1}, \dots , x_{\sigma d})$. 
The following theorem was proved by Gerstenhaber~\cite{Gerst1960}  for binary algebras and generalized  to the general case $d\ge 2$ by Umirbaev~\cite[Lemma 9]{Umirbaev}. 

\begin{Theorem}[Gerstenhaber, Umirbaev]
Each Gerstenhaber nil algebra is Engel. 
\end{Theorem}
In the binary case Gerstenhaber proved the following  estimate: if $A$ is Gerstenhaber $p$-nil, then it is $n$-Engel for  $n = 2p-3$~\cite[Theorem~1]{Gerst1960}.  For a generalization of this estimate to the case of $d$-linear algebras in a more general setting, see Theorem~\ref{th:main} below.

Another version of nilpotence appears in an equivalent formulation of the Jacobian problem.
 Let $T^{mult}_q (x_1, \dots , x_q)$ be the sum of all multilinear multiple compositions of $\mu$ with $q$ arguments, and let $T_q(x) = \frac{1}{q!}T_q^{mult}(x,x, \dots, x)$. 
In other words, 
$$
T_1(x) = x \mbox{ and }T_q (x) = \sum_{i_1 +\dots + i_d = q}\mu(T_{i_1}, \dots, T_{i_d}) 
\mbox{ for }q>1. 
$$ 

Let us call an algebra $A$ {\em Yagzhev nil} (of nilindex $p$) if the identities $T_q(x) \equiv 0$ hold in $A$ for all $q\ge p$. Such algebras are called Yagzhev, or weakly nilpotent  in~\cite{Belov_etc}, and weakly nil in~\cite{dots}.  Note that Yagzhev nil algebras need not be Gerstenhaber nil.  Indeed, 
Gorni and Zampieri~\cite{gz} have shown that there exists a 4-dimensional 3-linear algebra which is Yagzhev nil and Engel but not Gerstenhaber nil (their example is induced by a polynomial automorphism due to van den Essen).  However, in the important case of binary algebras we do not know whether these two nil properties are equivalent? At least, a straightforward calculation with linearized identities shows that binary Yagzhev nil  algebras of nilindex $4$ are Gerstenhaber nil of nilindex $6$.

Yagzhev nil algebras  appear in his reformulation of the famous Jacobian conjecture, see~\cite{Belov_etc, Umirbaev} and references therein.  Recall that the conjecture states that a complex polynomial endomorphism with constant Jacobian determinant has a polynomial inverse. 
 It is equivalent to the following statement (\cite{Yagzhev}; see also \cite{bass}):
 
 \begin{Conjecture}[Jacobian conjecture for homogeneous mapping]
 	\label{conj:Jacobian_hom}
 	Suppose that $F: \Complex^n \to \Complex^n$ is a polynomial automorphism of the form 
 	$F = \Identity - H$, where $H$ is a homogeneous automorphism of degree $d\ge 2$.
 	If the determinant $j_F$ of the Jacobian matrix $J_F = (\partial F(X)/\partial X)$ of the map $F$ is equal to the constant 1, then the map $F$ has a polynomial inverse. 
 \end{Conjecture}
 
 The conjecture is known to be true for $d=2$~\cite{Wang}, and it is sufficient to prove it for $d=3$~(\cite{Yagzhev,bass}). 
 
 It follows from the results by Yagzhev~\cite{Yagzhev2000} (based on an earlier work by Dru\. zkowski and Rusek~\cite{dr}) 
that the Jacobian conjecture for homogeneous mapping of degree $d$ for polynomials of $n$ variables  is equivalent to the following:
 
 \begin{Conjecture}[A Jacobian conjecture in Yagzhev form]
 	\label{conj:Yagzhev}
 	Suppose that $A$ is a complex $n$-dimensional  $d$-linear algebra. If $A$ is Engel, 
then it is Yagzhev nil.  
 \end{Conjecture}
 In the notation of Conjecture~\ref{conj:Jacobian_hom}, here $\mu$ is the complete linearization of $H$ so that $ H(X) =\mu(X, \dots, X)$.  


 The easy reverse implication of the Jacobian conjecture (stating that the Jacobian determinant of a polynomial automorphism is a nonzero constant) is  well known as the  Jacobian theorem. Being reformulated in the Yagzhev terms, it is a refinement of the Gerstenhaber--Umirbaev theorem for finite-dimensional algebras.

 \begin{Proposition}[Jacobian theorem]
 	\label{prop:Easy_part_of_JC}
 	Suppose that a  complex $d$-linear algebra $A$ is finite-dimensional.
If $A$ is Yagzhev nil, then it is  Engel. 
 \end{Proposition}
 
The aim of this note is to prove the same result without the assumption that $A$ is finite-dimensional. We regard this as an infinite-dimensional version of the Jacobian theorem.

 \begin{Theorem}[Jacobian theorem for an infinite set of variables]
 	\label{th:main}
	Each Yagzhev nil algebra $A$ is  Engel. 
If the Yagzhev nilindex of $A$ is $p$, 
then $A$ is $n$-Engel
for $n = d\left[ \frac{p-2}{d-1}\right]  +1$. 
 \end{Theorem}
Note that the assumption of the second implication  can be replaced by the following: the identities $T_q(x) \equiv 0$ for all $p \le q \le d(p-1)+1$ (cf.~\cite[Prop.~3]{dots}). 

 Dotsenko has asked whether, in the binary case $d=2$, the identity $T_q(x) \equiv 0$  implies the $n$-Engel identity  for some $n$ \cite[Question 4]{dots}?
Theorem~\ref{th:main} gives a partial answer to this question. 
Indeed,  if the identities $T_q(x) \equiv 0$ hold for all $p \le q \le 2p-1$, the algebra is $n$--Engel for $n = 2p-3$.  Moreover, one can show that if the identities $T_4(x) \equiv T_5(x)  \equiv 0$ hold in a binary algebra $A$, then it is Gerstenhaber nil  of nilindex 6 (hence, Yagzhev nil of nilindex 4); therefore, it is 5-Engel (cf.~\cite[Question~3]{dots}).

\medskip

{\it Proof of Theorem~\ref{th:main}. }\ 
 Let $q_0$ be the minimal  number such that the identities $T_q(x) =0$ hold in $A$ for all $q\ge q_0$. Since by definition $T_q(x) = 0$ for $q \notin (d-1)\Z +1$, we have
$q_0 = (d-1)N+2$, where the integer $N$ satisfies $N \le \left[ \frac{p-2}{d-1}\right]$.

 We may assume that all relations of $A$ are consequences of the identities $T_q(x) \equiv 0$ for all $q\ge p$, so that $A$ is a free algebra of some variety of multilinear  algebras. In particular, we assume that $A$ is graded. Moreover, we will assume that the set of free  generators of $A$ consists of at least two elements.
 
 The operator $g:A\to A$ defined by $g(x) = x - \mu (x, \dots ,x)$ is invertible; 
 the  inverse is given by $\gamma(y) := g^{-1}(y) = \sum_{j\ge 1} T_j(y)$~\cite{dr} (where $T_j(y)=0$ for $j\ge q_0$).
 We obtain the identities
 $$
 \gamma (g(x)) = \sum_{j\ge 1} T_j(g(x)) = x
 \mbox{ and }  g(\gamma(y)) = \gamma(y)-\mu (\gamma(y), \dots ,\gamma(y)) = y.
 $$
Replacing $x$ by $x+z$, $y$ by $y+t$, and collecting all terms which are linear in $z$ and $t$, we obtain partial linearizations of the above identities
 $$
 \mathrm{d}\gamma (g(x), \mathrm{d}g(x,z)) = z\mbox{ and }
 \mathrm{d}g (\gamma(y), \mathrm{d}\gamma(y,t)) = t,
 $$
 where 
$$\mathrm{d}\gamma(y,t) = \sum_{n= 1}^{q_0-1} \frac{1}{(j-1)!}T_j^{mult}(y,\dots, y,t)
\mbox{ and }
\mathrm{d}g(x,z) = z - d \AdOper_x(z)
$$ are the corresponding partial linearizations of $\gamma$ and $g$. 
 So, the linear operator $\mathrm{d}g_x: z \mapsto \mathrm{d}g(x,z)$
 (the ``Jacobian'') is invertible; the inverse is given by 
 the operator $t \mapsto \mathrm{d}\gamma(g(x),t)$.
 
 On the other hand, 
the linear operator $\mathrm{d}g_x$ can be extended to the 
completion $\widehat A$ of the graded algebra $A$ 
by the same formula $\widehat{\mathrm{d}g_x}: z\mapsto z - d \mu(x, \dots, x,z)$.
It admits an inverse  
 $\widehat{\mathrm{d}g_x}^{-1}: t \mapsto t+ \sum_{i\ge 1} d^i \AdOper_x^i(t)$. Since 
 the restriction of the operator $\widehat{\mathrm{d}g_x}$ to the subset $A\subset \widehat A$ is a bijection $\mathrm{d}g_x:A\to A$, we have $\widehat{\mathrm{d}g_x}^{-1}(A)\subset A$. So, the higher homogeneous components $d^i \AdOper_x^i(t)$ with $i>>0$ of the element $\widehat{\mathrm{d}g_x}^{-1}(t) \in \widehat A $   vanish for $x, t\in A$. 
 It follows that for each pair of generators $x$ and $t$ of  $A$ we have $\AdOper_x^i(t) =0$
 for some $i$. Since $A$ is a free algebra of some variety, this equality is an identity in $A$.

Let $n$ be the smallest number such that the Engel identity $\AdOper_x^n(t) =0$ holds. 
We have an equality
$$
\widehat{\mathrm{d}g_x}^{-1}(t) = {\mathrm{d}g_x}^{-1}(t)
$$
in $A$ for the generators $x,t$ of $A$. Since the algebra is graded, this equality implies the equality of corresponding homogeneous components. On the left-hand side  
$\widehat{\mathrm{d}g_x}^{-1}(t) =  t+ \sum_{i\ge 1} d^i \AdOper_x^i(t)$,
the highest nonzero component is the one with $i=n-1$; its degree is $(d-1)(n-1)+1$.
On the right-hand side  
$${\mathrm{d}g_x}^{-1}(t) = \mathrm{d}\gamma(g(x),t) = 
\sum_{j= 1}^{q_0-1} \frac{1}{(j-1)!}T_j^{mult}(g(x),\dots, g(x),t),
$$
the degrees of the nonzero homogeneous components
do not exceed the number $d(q_0-2) +1 = d(d-1)N+1$. Therefore, we obtain the inequality
$$
(d-1)(n-1)+1 \le d(d-1)N+1.
$$
Thus,   $n\le dN +1 \le d \left[ \frac{p-2}{d-1}\right]+1$.
 \ep\\
 
The converse of Theorem~\ref{th:main}  (whether an arbitrary Engel algebra is Yagzhev nil?) is a challenging problem which generalizes the Jacobian conjecture to the case of infinite number of variables.  For binary algebras, it is stated in~\cite{Belov_etc} as the Generalized Jacobian conjecture for quadratic mappings.  This last conjecture holds for power-associative algebras and for those satisfying the identity $(x^2)^2\equiv 0$~\cite[Sec.~8]{pz} as well as for 3-Engel algebras~\cite{singer}. Note that in all these cases the algebras turn out to be not only Yagzhev nil but also Gerstenhaber nil.  

\medskip
 {\large \bf Acknowledgments}

I am grateful to Vladimir Dotsenko and Fouad Zitan for stimulating discussions.  
 This work was supported by the Russian Science Foundation, grant RSF 24-21-00341.

 \end{document}